Sally can paint a room in 7 hours while it takes Steve 9 hours to paint the same room. How long would it take them to paint the room if they worked​ together?
Answer by Boreal(15235)

in 1 hour, Sally paints 1/7 of the room and Steve 1/9 of the room.
In x hours, working together, they paint
(x/7)+(x/9)=1 full room.
multiply by the LCD of 63
9x+7x=63
16x=63
x=(63/16) hours or just under 4 hours (3.9375 hours)